A chemist wishes to mix a 10% acid solution with a 60% acid solution to obtain 300 quarts of a 24% acid solution. How many quarts of each solution should be mixed together?
Answer by TimothyLamb(4379)

x = quarts of 10% solution
y = quarts of 60% solution
---
10x + 60y = 24*300
x + y = 300
---
put the system of linear equations into standard form
---
10x + 60y = 7200
x + y = 300
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x = quarts of 10% solution = 216
y = quarts of 60% solution = 84
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php